OpenStudy (egbeach):

Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°

OpenStudy (amoodarya):

 \[\tan \theta =-1 \rightarrow \theta=2\pi-\frac{\pi}{4}\\r=\sqrt{x^2+y^2}=\sqrt{3^2+(-3)^2}=\sqrt{18}\]

OpenStudy (egbeach):

its asking for the coordinates though